OpenStudy (liddoislander):

The attendance at a football game last year was 5,600. This year the attendance was 4,560. What was the percent of decrease in attendance, to the nearest percent? A. 19% B. 23% C. 81% D. 104%

OpenStudy (shamil98):

Use the formula \[\huge \frac{ part }{whole} = \frac{ \text{%} }{ 100 }\] \[\huge \frac{ 4560 }{ 5600 } = \frac{ x }{ 100 }\]

OpenStudy (shamil98):

\[\huge 456000 = 5600x\] \[\huge 81.4 = x\] so 4560 is 81% of 5600. so it was decreased by 19%.

OpenStudy (liddoislander):

thank you . you were so helpful !! Much alofa (love ) to you !!

OpenStudy (shamil98):

Glad to be of help.